Information for Math 254-7 Differential Equations Fall, 1996

Instructor: Doug Ulmer  Class: Tuesdays 12:30-1:45, PAS 304
Office: Math 716  Thursdays 12:30-1:45, Math 102
Phone: 621-6861  Office Hours: Wed 1-2 in the Tutoring Room

Thu 2-3:30 in my office

   E-mail: ulmer@math.arizona.edu
  Website: http://www.math.arizona.edu/people/~ulmer/

              teaching/DiffF96/DiffF96.html

The course: This course is an introduction to differential equations. We will learn how to solve some equations explicitly when this is possible, how to decide when it might be possible, and how to find key properties of solutions even when they cannot be found explcitly in terms of simple functions. The emphasis throughout will be on understanding the equations and their solutions from the analytical, graphical, and numerical perspectives. “Cookbook” recipes for solutions will not be important in this course.

Text and Technology: The text is “Exploring Differential Equations via Graphics and Data” by Lomen and Lovelock. There is a student solutions manual which should be available in the bookstore in about two weeks. You will need to buy and learn how to use a graphing calculator (for example a TI-82). We will also be using the U of A’s public domain math software both in class (on Thursdays in Math 102) and outside of class (in public access labs or at home).

Homework and Quizes: There will be two sorts of homework, “Exercises” and “Problems”. Exercises will be assigned daily, are required, but will not be collected. Most exercises will have solutions in the student solutions manual so you can check your work; I will distribute solutions for the rest. There will be occasional short quizes in class based on homework exercises. The first quiz is on Thursday, 8/29 .

Problems will be assigned and graded in groups. These will be longer and more difficult and will require you to meet with your group, work together on solving the problem and on writing up the solution. All members of the group will receive the same credit for the problem. I will assign groups (of 3-4 students each) and the groups will change from time to time. This form of work may be unusual for you (especially in a math class!) but it is much more like the real world than traditional homework.

Exams: There will be in class exams on Thursday 9/12, Tuesday 10/22, and Thursday 12/5. The final exam is Tuesday, 12/17 from 11 to 1 in PAS 304.

Grades: Grades will be based on homework and quizes, the three hour exams and the final, in the following proportions:

Homework and quizes  25%
Hour exams  15% each
Final exam  30%

Course grades will be determined by a straight percentage scale: 90-100% is an A, 80-90% is a B, 70-80% is a C, and 60-70% is a D.

Policy on Withdrawals and make-ups: If you withdraw on or before 10/30, you will receive a W if you are passing the course at the time of withdrawal (defined as having completed 60% of the work due at that time).

Late homework is unfair to everyone and is strongly discouraged; depending on cicumstances, it may not earn full credit. Make up exams will be allowed only with the most compelling need. Don’t even think about missing an exam without making prior arrangements.

Tips for success:

1) Show up for class, ready to work. Experience shows that very few students can succeed in this course without coming to class every day.

2) Do the assigned reading and homework exercises before class. Class time will be spent working interesting examples and dealing with more difficult points, not repeating what is in the book.

3) Don’t get behind! This class moves fast it will be very difficult to catch up if you fall behind. Try to get some help (in office hours, in the tutoring room, or from fellow students) at the first sign of trouble.